OpenStudy (jenniferjuice):

Solve: 5(x + 8) = 7(x − 2)

OpenStudy (uri):

5(x + 8) = 7(x − 2) 5x+40 =7x -14 40+14 =7x -5x 54=2x x=54/2 x=27.Ans :D

OpenStudy (anonymous):

I agree their math checks out.

OpenStudy (jenniferjuice):

thankyou